CCD Imaging Spectrometer
George C. Marshall Space Flight Center
National Aeronautics and Space Administration
Marshall Space Flight Center, AL 35812
Center for Space Research
Massachusetts Institute of Technology
Cambridge, MA 02139
This report covers the period January 1996.
The monthly status meeting for ACIS was held on the morning of January 17, at MIT. The review was attended by Max Rosenthal and Jack Loose from MSFC. Gordon Garmire was unable to attend due to illness and George Ricker was traveling on NASA business. This monthly review was followed by a MSFC requested review of the overall ACIS Verification Testing program in the afternoon. Ed Sedivy from LMA attended both meetings and presented the LMA status and verification plans for the PSMC. The main reason for this Verification review was that, due to schedule problems, if the thermal vacuum testing of the ACIS instrument is moved to Lincoln Lab, it would have to be done without the flight PSMC. Then the only totally integrated environmental test of ACIS prior to delivery would be EMI/EMC at LMA, Denver. By January 18, MSFC gave us approval to proceed with these plans and meetings were immediately set up with Lincoln Lab to begin planning for the ACIS thermal vacuum testing in August of this year. The next monthly review is scheduled for March 6. MSFC has requested that the review be at LMA, Denver, since the critical path for ACIS delivery is now gated by the delivery of the flight PSMC.
The monthly Technical Interface Meeting with LMA was held on January 25 and 26 in Denver. Five members of the MIT/ACIS staff traveled to Denver for this meeting. Due to the numerous topics to be discussed, up to three splinter meetings were running in parallel most of the time. From a programmatic point of view, the main points are that an LMA reorganization has been completed and Lloyd Oldham now reports to Ed Euler; a solution to the protracted Detector Assembly alignment process has been proposed by LMA and is under review; and all LMA hardware appears to be holding schedule but the PSMC is very tight. Due to the monthly review date and location, the next TIM is also scheduled at LMA for March 4 and 5. Tentative plans also call for the MIT subcontract officer to travel to LMA for March 7 so that face-to-face negotiations can be conducted for Change Order #43 (and #25 if necessary).
MIT participated in the AXAF project level telecons on January 9, 16, and 30. There was no AXAF telecon on January 2 due to the shutdown of the Federal Government, and ACIS did not participate in the telecon on January 23 due to an ACIS Science Team meeting conducted at MIT on that date.
MIT participated in the ACIS bi-weekly status review on January 10. The ACIS telecon normally scheduled for January 24 was canceled due to the monthly status review the week before.
During the month of January, the ACIS schedule continued to be a problem. The schedule held but only with much re-planning and juggling by both MIT and LMA. LMA entered the equation with a late delivery of the PSMC (due to the change in requirements by MIT) and the Vent Valve Assembly (due to weld problems on the flight tube assembly). By the end of the month, LMA had found a new welding company which has (at least promised) a faster delivery and the Vent Valve Assembly moved back off the critical path. The late delivery of the PSMC has caused a re-planning of the Verification Plan, with separate thermal vacuum tests for the DPS and the PSMC. MSFC approved this plan on January 18, so the schedule held.
Schedule issues remain in the area of Detector Assembly alignment time (LMA estimates about 33 days vs. a schedule allowance of 19 days), and the impact of modifying the flight Collimator to support the Proton Shield components provided by MIT. [Due to confusion introduced by the MSFC/AXAF Science Office report that claimed no proton shielding was necessary, this area of the ACIS design was ignored, and now the fully completed flight Collimator must be modified with Proton Shield mounting holes].
As of the end of January, we are still showing seven weeks of slack in the overall ACIS schedule which results in the delivery of the ACIS experiment to BASD on December 15. However, only three weeks of this slack remain after the scheduled delivery of the PSMC on October 8 and are available to resolve LMA schedule problems.
The issue of single point failures on the Analog Boards was resolved. The PSMC has redundant power supplies (six regulated voltages) for the DEA. These redundant voltages were diode-ORed on each Analog Board. This resulted in six single-point failures on each of the ten Analog Boards in the DEA. Although the probability of a short at any one of these points is very low, the fact that there are sixty such points was very worrisome. A fair amount of effort was expended in December and early January to try to develop an alternative approach to dealing with the DEA power distribution. A meeting was held on January 5 and the decision was made to resolve the problem with the introduction of about 15 double pole latching relays on the DEA Interface Board. The Analog Boards are paired and each pair receives its regulated voltages through one pole of one of the relays. Hence, a failure is now restricted to two boards, not the entire DEA.
As discussed above, the ACIS project is just holding from a schedule point of view. Sufficient front illuminated (FI) CCDs are on hand to populate the flight Focal Plane (but the BI CCDs are having trouble). Five Analog Board test sets have been delivered to the science team and the calibration of the flight CCDs has begun. Actual flight hardware is beginning to arrive; the flight OBF frames were received and sent to LMA for cleaning, the dip-brazed connector panel for the DEA was received and accepted.
The main technical problems of the past few months seem to be under control. The weight and power allocations were maintained through January.
During January, several personnel changes were made to the ACIS staff at MIT. The MIT project decided to hire a contractor software test engineer (for about seven months) in order to keep the flight software effort on track, and he started work in January. As indicated last month, David Voutour, the design engineer for the Back End Processor of the DPA, left MIT on January 19. Due to the short term nature of this position (the job was scheduled to end in May), the ACIS project management has decided not to replace David. Dorothy Gordon (FEP design engineer) has agreed to take on the responsibility for the BEP design. In order to provide Dorothy with some help, we will also augment the SOW of Jim Littlefield - a contractor already working on the testing of the FEP and BEP.
Nine devices, all front illuminated, were received in January. Seven of these are flight candidates; the other two are engineering test devices. All flight candidates passed the screening test. There are now 20 devices at CSR which have passed screening tests, and are available for use in the flight Focal Plane. They are itemized below:
|Device||(FI/BI)||Date Rec'd||Date Screened
|w147c3||BI||10 Aug||11 Aug|
|w129c2||FI||15 Sep||15 Sep||Shows "D noise"|
|w87c4||FI||3 Nov||3 Nov|
|w78c1||FI||7 Nov||8 Nov|
|w97c1||FI||17 Nov||17 Nov||Shows "D noise"|
|w157c1||FI||24 Nov||24 Nov|
|w157c2||FI||24 Nov||27 Nov||Shows "D noise"|
|w158c2||FI||4 Dec||4 Dec||Shows "D noise"|
|w158c4||FI||4 Dec||5 Dec|
|w159c1||FI||14 Dec||15 Dec||Marg. spectral resol. at OK|
|w163c3||FI||14 Dec||19 Dec||Shows "D noise"|
|w168c2||FI||19 Dec||22 Dec|
|w168c4||FI||19 Dec||26 Dec|
|w163c1||FI||10 Jan||19 Jan|
|w182c4||FI||10 Jan||22 Jan|
|w185c3||FI||10 Jan||23 Jan|
|w189c2||FI||17 Jan||24 Jan|
|w186c4||FI||17 Jan||26 Jan|
|w190c1||FI||23 Jan||1 Feb|
Performance data for these devices is available on the ACIS web site.
The calibration electronics tests (Rev. B DEA boards) were received in January. After an intensive testing effort, a number of problems with the boards were identified and solved. In order to obtain the required linear ity, it was necessary to modify the timing of the analog-to-digital conversion. Kissel discovered that reduction of the input (CCD JFET load) resistance improved noise performance substantially. With this change, noise levels below 2 electrons, RMS, were obtained with flight detectors for the first time. At the end of January, we were awaiting modifications of the calibration test sets to implement these changes. A number of other questions and issues regarding operation of the DEA board design remain; most of these issues affect the flight DEA board design more than the utility of the calibration test sets.
High-energy calibration of three reference CCD detectors, with respect to the Si(Li) detector, was completed.
Work continued to optimize the final CCD operating clock levels.
A science team meeting was held on 23 January at MIT. The principal topic of the meeting was the placement of back-illuminated CCDs in the ACIS Focal Plane. Minutes have been distributed by the Principal Investigator.
Two graduate students from Pennsylvania State University, Kaori Nishikida and Catherine Grant, took up residence at CSR. They began assisting in the calibration data acquisition and analysis. Beverly Ferguson continued her support of data acquisition activities and procedure preparation.
The ACIS Contamination Monitor (Door source) engineering unit parts are out for gold plating. The drawing package for the flight unit has been reviewed.
The External Calibration Sources are on order. The existing drawing package was checked prior to release. Minor changes to the collimation of the beam are required. Materials are being ordered.
Tooling is being prepared to permit leak checking of the Back Plate assemblies.
The design of the Proton Shield is proceeding. Mounting holes have been defined so that LMA can include them in the flight Collimator.
Detailed drawings of the Structure are released and the large parts are being machined. The DEA to DPA cable routing has been included in the mock-up. Design of the brackets for the DA to DEA cables is nearly complete. Reinforcement of the Flexure to Support Structure joint is planned for the flight unit. We are concerned with Ball's decision to not fully support the feet of the Flexures. This may require stiffening of the Flexures.
The engineering unit Focal Plane has been received from Lincoln Labs. Assembly Work Orders are being prepared.
All parts for the engineering unit have been received, however, the mechanical parts are out for plating.
Flight frames are being cleaned at LMA.
The lifting fixture for the Support Structure, DEA, and DPA is being built for the mock-up and the engineering unit.
The Detector Electronics Assembly engineering unit enclosure and aluminum printed wiring board frames have been assembled. No fit problems have been encountered. The assembly is now mounted to the support structure mock up to fit the CCD and interface cables. Dip brazed Z panels for the Digital Processing Assembly have been received and evaluated. One panel has been designated for flight.
Revision C of the DEA CCA in Copper/Invar/Copper has been received and is in assembly. Minor packaging modifications to the TO-5's are in process. The Preparatory Board has been eliminated from the DPA. The receivers and memory have been relocated to the BEP and FEP.
Consulted with MSFC/TRW to obtain approval to delete redundant ACIS RCTU thermistors from the baseline design.
Participated in thermal telecons on 1/8, 1/9, 1/16, 1/23, and 1/30 as a member of ISIM Thermal Tiger Team to resolve ISIM thermal issues related to 173W ACIS.
Negotiated locations of ACIS operational and survival heaters with members of ISIM Thermal Tiger Team. The heaters were moved from the -Z wall of the SS to the +Y and -Y walls of the SS. The new locations reduce the overall thermal gradient from the heaters to the coldest DEA/DPA box walls and preclude the survival heaters turning on during a 2-hour eclipse.
Completed the ACIS transient thermal model.
Revised DPA-FEP thermal model to reflect updated design.
Analyzed the DEA I/O and TC board hottest electronic components.
Visited Lincoln Labs to discuss the ACIS TV test planned for July 1996.
The DEA engineering unit subsystems are performing exceptionally well. Five full systems have been tested, debugged and delivered to the scientists for flight CCD calibration activity.
Effort continues on the design and development of the 11th DEA card which serves as the system I/O interface and Focal Plane heating control. Command and data interface issues are being resolved and will impact flight designs.
The FEP engineering board layout is ongoing: parts placement is almost completed; routing is beginning. In parallel, the FEP bread boards are still being run in diagnostic mode, with incremental Actel fixes being generated, as required.
David Voutour has left CSR, and it was decided that the best course would be to redesign the BEP from scratch. In the meantime, the breadboard BEP (with software incorporating various BEP hardware bug workarounds) is still usable as a preliminary flight software development platform.
The BEP engineering board design has begun. A preliminary placement revealed that it was feasible to fit the eight flight EEPROMs on the BEP, making it possible to rid the system of the PREP board. This change was accomplished, and incorporated in the FEP engineering board layout. (It involves receiving DEA pixel data differentially on the FEP, rather than the previous configuration of single-ended reception.) The elimination of the PREP board not only decreases the total weight of the DPA, but also simplifies the system design, testing and manufacturing.
Three "Frame Buffer" units have been fabricated. Functional testing of two of the units is now complete. Although a SPARCstation has been configured with a Direct Memory Access interface, it has yet to be tested with a Frame Buffer.
The ACIS thermal vacuum test configuration plan and design of the MIT supplied harnesses is in progress.
Developed a preliminary version of filterServer and filterClient software and tested them. Began developing software to read ACIS-format telemetry packets from an input stream. Met with LMA personnel and discussed changes to the GSE Contract End Item Specification.
Held 2 telecons with TRW and MSFC to discuss the issue of ACIS commanding. TRW is concerned that dropped command words will place ACIS into an undetermined state. This may result in changes to BEP command management, but no decision has yet been taken. A description of ACIS command and telemetry formats has been made available on MIT's ACIS page of the World Wide Web ("http://acis.mit.edu/ipcl/").
Three (out of 18) action items remain open.
Conducted an ECO review of Huffman Compression classes.
Continued testing the BEP serial telemetry interface to characterize various firmware problems. The BEP-DEA interface is largely debugged, and the execution time for various functions carefully measured.
The pre-alpha release is ready for testing by Q/A. Some preliminary tests have been run using the BEP simulator, but serious testing will have to await the completion of BEP hardware testing.
All BEP and FEP flight software modules continue to be subjected to unit and coverage tests.
A BEP software simulator has been developed and tested. Work has begun to join it to the FEP simulator in order to test the full instrument software. Work continues on the suite of software test tools. The ACIS command builder is complete, pixel loaders are complete and the telemetry parser is 75% complete. Software has been developed to create test telemetry streams from realistic CCD data. These will be used to exercise the test tools and will also be sent to PSU and ASC following their recent request.
Alan Braver has joined the software Q/A effort in the area of test script preparation and documentation. His work will be directed through Q/A management and he will attend all future flight software meetings.
No Alerts were received from NASA/MSFC over the report period. MSFC Alert #6889, which includes the MS27505E17F8PA (A1J1) and MS27505E17F8PB (A1J2), were reported last month. These two (2) connectors were purchased by MIT and supplied to Ball, who is building the interface cables. Both TRW and Ball were informed of the Alert and the MIT shipment to Ball. The two (2) connector types have been reordered with an expected ship date of 4/30/96.
Generated five (5) additional DEA kits (SN 9 - 14). All were issued.
Prepared a job description, solicited resumes, interviewed perspective applicants and hired a temporary stock/data entry clerk to assist in the operation of the bonded stock room.
Seven component lots were hand carried to Associated Testing Laboratories, Inc., Burlington, MA for multi-stage part testing (PIND, X-ray, DPA).
Seven parts were lead formed and tinned at Lockheed Sanders, Nashua, NH.
Flight resistors from Mini-Systems were received and inspected. Seven lots were mismarked and returned to Mini-Systems for re-marking.
Final source inspection was done at Chip Supply on the OP220 (37-02307).
Final source inspection was done at Harris on the HS9-26C31RH-8, HS1-26C31RH-8, and ACTS244KMSH.
Final source inspection is planned for the week of 2/12/96 at SEI on the 36-02306 (EEPROM).
Final source inspection is planned for the week of 2/12/96 at Q-Tech on the 36-02311 (38.4 MHZ Crystal Oscillator).
Nitrogen supplies have been delivered and are ready for installation. The Nitrogen will be used for purging bags for shipping, as well as blowing of particles and drying material before admittance into the cleanrooms.
A fifteen foot cable was made for use on the TQCM at Lincoln.
A change in bagging material location was made. Nylon 6, clean to level 50A, will now be used as an inner bag and the Aclar 22C, clean to level 50A, will be used as an outer bag.
Sent four parts out to MSFC for 1443 testing:
Deviation request numbers 36-002, 36-003, 36-004, and 36-006, regarding Lockheed Martin Astronautics (LMA) procedures and processes, were approved with comment.
Prepared SCD 36-02312 for a Polypropylene Capacitor. This will be submitted to MSFC as NSPAR 36-024.
No new NSPARs were submitted for approval during the report period. No NSPAR approvals are outstanding.
|36-002||A to D Converter
|36-003||CA Memory Module
|36-004||FB Memory Module
|36-005||Programmable Supply current
Op Amp 36-02304
Only Memory 36-02306
|36-008||Electrical Connectors, PCB
Mount SND Type
|36-009||Electrical Connectors, PCB
Mount KA Type
|36-012||Junction Field Effect Transistor
|36-013||Dual Surface Mount Diode
|36-014||Dual Operational Amplifier
|36-015||8000 Gate Anti-fuse Field
Programmable Gate Array
|36-017A||Charge Coupled Device
|36-018||Microcircuit, Octal Buffer
|36-019||Microcircuit, Octal Bus
Transceiver (Harris HCS245)
Flip-Flop (Harris HCS374)
Line Driver (Harris HS26C31)
Differential Line Receiver
Q-Tech part type
MIT has encouraged Lockheed Martin Astronautics (LMA) to submit required NSPARs to MIT. These are in process at LMA.
Radiation testing has been completed at Space Electronics, Inc. (SEI) on fourteen (14) device types. Results of these tests are listed below.
|Manufacturer||Part Number||Radiation Test
|Analog Devices||DAC8800BR/883||<2K Rads|
|Com Linear||CLC505A8D||>100K Rads|
|Harris (Chip Supply)||36-02305 (CA 3080)||<100K Rads|
|Analog Devices||OP220AJ/883 (TO-5 can)||8K Rads|
|Texas Instruments||TL082/883B||>100K Rads|
|Analog Devices||REF43BZ/883||>200 K Rads|
|Chip Supply||OP-220 (DIP)
Devices which have not passed 100K Rads of Cobalt 60 testing will be shielded or design work-arounds will be implemented. Eight (8) more device types are planned for radiation testing. These are listed above without results.
Five Teledyne Frame Buffer Modules were forwarded to SEI, San Diego, CA for radiation testing.
Three lots of components were forwarded to Interpoint, Redmond, WA for radiation shielding:
LMA has been instructed to remove the Ni bottle from the MGSE while it is at KSC.
There has been no activity on the Performance Assurance and Safety (PAS) Plan. The PAS Plan in effect is revision B.
The flight OBF frames (8 ea.) and shipping containers (6 ea.) were sent to LMA for cleaning, prior to shipment to Luxel for filter installation.
Fifteen OWS holders were cleaned at Lincoln Labs.
The cleaning list for the DEA, DPA, Support Structure, and GSE has been generated and distributed for comment prior to updating the database.
The vacuum chamber for CCD calibration testing was cleaned and certified to level A.
Continuous efforts are being developed for particulate sampling using the tape test method per ASTM F 312-69, and NVR sampling using solvent wipe per MIL-STD-1246 using the facilities in the HETG cleanroom.
Developed the plans necessary to build and clean flight cable assemblies. Material was sent to Lincoln Labs for preconditioning and cleaning.
Continue to develop an overall procedure which includes dressing, behavior, material handling, work area preparation, flight material storage and bagging in cleanrooms for contamination control.
Continuous monitoring of the CCD calibration testing vacuum chamber is under way.
One inch optical witness samples have been installed onto each flight CCD.
Hydrocarbon levels are being maintained to acceptable levels.
Participated in Software ECO CCB: ECO 481. Huffman Table Class.
Continued developing memory access test procedure.
Participated in test tools review session.
Prepared a job description, solicited resumes, interviewed perspective applicants and hired a temporary engineer to assist in the testing and validation of the ACIS software.
During this month six flight candidate FI chips were shipped to campus. All six of the FI flight candidates were considered acceptable, raising the total number of FI candidates at CSR to twelve, which is more than adequate to populate the first flight paddle. There are sufficient FI candidates at various stages of the packaging process to yield a total of forty FI flight devices. The gating item for packaging is now considered to be flex prints, although an adequate supply is due for delivery at the beginning of January.
The emphasis in packaging will now be switched to delivering BI parts to determine how many can be incorporated in the first flight paddle. Two devices are in the packaging sequence now and will be ready for delivery early in January. There are six more BI candidates besides those produced in BI Lot 4 (which suffer from high dark current and high CTI), suggesting that a total of four BI chips could be available for use in the first flight paddle, or as backups.
Eddy current testing of the flight back-up structures has been successfully completed. Locating pins and threaded inserts have been installed and locations verified. These parts have been sent out for gold plating which will complete the fabrication process.
All components required for assembly of the flight unit are in hand except for the heater elements used for temperature control and contamination bake-out. These heaters are expected during the first week in February.
Additional components to assemble Detector Assemblies (alumina substrates and beryllium tees) have been ordered, to allow packaging of all available CCD devices. These components are expected to arrive long before existing supplies run out.
Metallization of the flexprint circuits for the low thermal emissivity coatings continues to proceed smoothly. All of the available flexprint circuits have been metallized successfully. Additional flexprint circuits are expected during the second half of January. These should arrive soon enough to prevent an interruption in Detector packaging.
During this month an update proposal was completed for Change Order 18 and fact finding was completed for Change Orders 39, 41 and 42. Negotiation of these Change Orders is scheduled for mid February.
The ACIS Program Technical Interchange Meeting (TIM) was held in Denver on January 25 & 26, 1996. This TIM centered around schedule coordination and working items with schedule concerns. Specific actions were assigned to reduce the schedule time required to accomplish the Focal Plane alignment verification, to recover from welding anomalies in the flight vent tubes and defining actions and approaches to meet the PSMC scheduled established by Change Order 43. The engineering unit Focal Plane was also packaged to enable the MIT personnel attending the TIM to hand-carry it back to MIT for post-test evaluation.
The Science Instrument Module (SIM) Thermal Tiger Team was supported this month through both attendance at the meeting in Boulder, CO and acceptance and response to several action items relating to optional mounting schemes for the PSMC. All requested actions have been completed and responses provided to the Team. Work on these actions has delayed scheduled completion of LMA thermal and stress analyses.
Major accomplishments for January included successful completion of the PSMC EU#1 performance tests following EMI/EMC and the thermal cycle tests. A program decision was made to keep the PSMC EU#1 at LMA as long as possible to help accelerate the checkout of the EU#2 PSMC in an attempt to resolve schedule creep concerns.
During January the reworked flight Vent Valve tube welding failed X-ray inspection and the LMA precision cleaning facility was shut to make safety improvements. These two unrelated anomalous events were resolved and workarounds were implemented. These corrective actions were made in time to prevent impact to the hardware need dates and, therefore, should have no impact on the flight hardware delivery schedule.
During the month the training and certification of LMA technicians and Quality Assurance inspectors for fabrication in compliance to NASA SPEC-NHB 5300.4 soldering was completed.
The anticipated Science Instrument Module (SIM) cabling routing data from Ball Aerospace was not provided in January. This data was requested by LMA in April 1995 and identified as needed no later than 17 January 1996 to support the LMA cable design and fabrication schedule. To attempt to accelerate development of this data, LMA prepared an alternative cable routing recommendation and provided it to Ball Aerospace. Work on ACIS flight cables has been delayed pending response to this recommendation and receipt of the cable routing data. Work around approaches are still being investigated to try to accommodate this late data delivery and avoid impacting the ACIS hardware delivery date.
Rework, recleaning and bakeout of the 3' x 5' chamber was completed this month. The chamber is continuing to be operated to assure acceptable cleanliness levels are achieved. The chamber will be certified in February before 1238 bake out activities begin.
Some EU #1 activities continued into this month. Additional thermal testing was performed to obtain information necessary for box thermal design verification. Significant effort was applied to incorporation of as-built redlines in preparation for delivery of the EU #1 drawing package. The table top review process was instrumental in identifying several areas that needed improvements on the EU #2 and flight versions. A successful functional test was performed with the EU #1 driving a StarSys Actuator.
Assembly of all the second engineering unit (EU#2) boards was completed. Individual PWB testing was started and will continue into next month. Changes that are driven by change orders and problems identified during internal reviews are being incorporated into EU #2. To aid in the efficient checkout of box level performance, a request was made to allow the EU #1 to remain in Denver as long as possible. This would allow integration and checkout of EU #2 boards into a known working box.
The PSMC Flight Equipment Specification was released, paving the way for release of all additional PSMC flight effectivity drawings. The flight magnetics drawing was released to get fabrication under way. There were no other flight drawings released during January. Parts procurement for flight build is proceeding as planned with no major problems to report.
Board and box level test tools received significant attention in getting modifications to match the changes that are occurring on the flight side. All additional required parts have been identified and ordered.
All test plans for radiation testing of electronic parts were completed and submitted for review. Work is continuing with our internal ra diation test lab to maximize commonality of data and EEE part test fixturing with other and prior EEE piece part testing to improve cost and schedule associated with this activity.
All work on flight Cable Harness design is, for all practical purposes, on hold pending cable routing definition from Ball Aerospace. Some limited work is progressing on 3D mockup fixturing for localized cable routing in the "Basement" (i.e. betwixt and between the DEA, DPA, DA, and Venting Subsystem). Very little more can be done in this arena without assuming risk, should the dictated routing paths from the "basement" to the PSMC require reroute.
Additional test harness definition is firming up for the new Lincoln Lab T-vac configuration and the integrated instrument EMC test at Denver, prior to instrument delivery.
All Detector Housing piece parts fabrication was completed in December and the small parts have now been precision cleaned. The bagging material contamination problem reported last month was resolved by selecting an alternate bagging material. The originally selected bagging material, ACLAR, was prone to static charge build up and attracted dust, lint, etc. Therefore, a new bagging material, which does not build up charge as easily, has been chosen. This has allowed the parts to be final bagged. Cleaning of the Collimator and Camera Body has been delayed pending MIT definition of a change to the Collimator to allow the attachment of addition Proton Shielding. The input which will define the locations of the Proton Shield mounting holes is expected in early February. As a result, the start of assembly of the Detector Housing and Collimator will be delayed approximately 4 weeks after these hole locations are defined. The four weeks is required to update and release the drawing, perform the machining, perform dye penetrant inspection (or another type in spection), and precision clean. A shear pin between the Camera Body and Collimator will also be incorporated into the flight Collimator and Camera to assure no relative movement during launch coincident with these changes.
Fabrication of TCS components is on schedule with no major problems to date. The warm radiator is in final machining, the cold radiator is in dye penetrant inspection, the warm and cold straps are in quality inspection, and the radiator standoffs are in final heat cure. Most of the other small details are complete and have been precision cleaned. Flight MLI fabrication will begin in February.
The PSU OBFs were received and inspected. Thermal cycling between -76 deg C and +46 deg C will commence in early February. The CAMSIM Focal Plane was removed and shipped to MIT for post-vibration/acoustic tests inspections. After inspection, the Focal Plane will be returned to LMA for the final OBF development test. The filters will then be installed into the CAMSIM and vibration tested in the X-axis with a 20 torr internal pressure and a 40 torr internal pressure for 60 seconds in each test. An optional 80 torr test will be included in the procedure in the event that additional data is desired. Previously the CAMSIM was only tested up to 40 torr in vibration test and 80 torr in acoustic test.
The completion of the stress analysis CDRL continues to be delayed as a result of more tasks assigned related to tiger team activities and action items. Heavy ACIS Tiger Team activities have required significant effort to support the PSMC base plate redesign. In addition, the delayed stress analysis now overlaps with the scheduled GSE drawing review and associated analysis which causes additional delays. These delays are not considered critical since no stress concerns have arisen and all components analyzed to date continue to show significant positive margins.
The PSMC piece part thermal analysis was updated in January. A complete spread sheet was generated for all the electronic parts in the box. This will allow the component temperatures to be automatically updated as the PSMC board and chassis temperatures change. The current predictions show some of the components at very near their maximum allowable junction temperature of 125 deg C for a +46 deg C PSMC side wall temperature. The high power axial lead diodes are up against their limits because they are very difficult to heat sink. However, if the current Ball prediction of +35 deg C (PSMC side walls) is maintained, all components will be within their derated temperature limits. As each of the boards are reviewed in table-top reviews, reference designators, piece part dissipations, and component locations will be modified to reflect the flight configuration.
The final failure analysis report from StarSys was received and forwarded to MIT. The test report for the life cycle test will be completed in early February.
The final reports will be released on the Door Actuator and Venting Subsystem life cycle tests once the final test report is received.
As a result of the vent tube welds failing X-ray inspection, a decision was made to procure new vent tubes from an alternate vendor. This decision was required because only two of the initial buy are still available, and if one of these was damaged during repair ACIS would not have a flight spare. Also, the new tubes can be ordered with all of the changes that have been made to the design incorporated with the new build. The procurement followed a set of vendor qualification tests which assured the new candidate vendor's processes produced a part that would pass flight quality inspections. The new vent tubes should be received by mid February and will be inspected at that time. This delivery date is consistent with hardware test and assembly need dates.
The VGSE mechanical drawings are nearly completed and will be released after review in early February. VGSE schematics and final as sembly drawings are progressing well and should be complete by the end of March.
Shipping container design and fabrication has been performed as required to meet the EU hardware shipping requirements. The design of the flight shipping containers will begin coincident with the definition of the SIM simulator design and shipping concept final definition. This is anticipated to begin in February.
The light shade analysis was completed in January. The results indicate that the internal vent tube will require non-reflective surfaces; i.e. black paint, in at least one of the elbows. This improves overall performance to the point where Earthshine is no longer a problem. The selected paint is Chemglaze Z306, which is on the MSFC-SPEC 1443 list as passed. This technical solution represents an absolute worst case analysis and, therefore, does not impact mission operations. Contamination effects from this are still under investigation.
Review of the Contamination Control Plan remains in process at LMA during this reporting period. A new release date is scheduled for 2/15/96 to accommodate incorporation of on-project comments.
The ACIS FMEA/CIL is being prepared at LMA. Collection of FMEA and CIL update information for ACIS DPS hardware is continuing to be developed. A DPS block diagram was created and circulated to the ACIS program for comments. Incorporation of the collected comments is scheduled for February 1996. During this reporting period LMA was requested to develop a reliability prediction for the PSMC to evaluate the effect of operating both PSMC to DPA power supplies simultaneously. This activity is being planned and will be scheduled to commence during the next reporting period.
EMI/EMC test report of the engineering unit PSMC test is being prepared during this reporting period. Preliminary results indicate that the PSMC met all requirements with margin, however, verification of this was clouded by emissions from the support equipment and the test set-up. The report is scheduled for publication during the next report ing period.
Results of the Spacecraft Grounding TIM were favorable. During this TIM TRW conducted testing of the ACIS to RCTU configuration that included a 3 box model of the spacecraft RCTU, ACIS DPA and ACIS PSMC. The results of the tests showed that a return wire between the DPA and RCTU was necessary to make the interface work. Excessive noise (EMI) at the receive end of circuits made the telemetry unread able by the RCTU. Future testing is planned to encompass the PSMC interfaces as soon as opto-isolators used in the PSMC could be provided to TRW. LMA provided these opto-isolators by Federal Express. Results of the remaining portion of testing will be reported next month.
Near the end of the Spacecraft Grounding TIM an issue was uncovered between the RCTU and PSMC. The issue is that the RCTU only accesses telemetry data from a single channel once during a major frame and the PSMC telemetry interface requires multiple accesses (4 times) of a single channel during a major frame. This did not represent the understanding that ACIS had of the interface. This item will be resolved prior to or at the Spacecraft CDR to be held next month.
Dialog continued on spacecraft charging concerns with TRW and ACIS throughout this reporting period. During this dialog a possible interference with the R.F. telemetry communications link caused by an electric field spectrum created from arc discharges was defined. This resulted from a TRW review that the spectrum caused arc discharges which considered the levels at 2 GHz too insignificant for the telemetry receiver. The interference level would add 0.1dB to the noise contribution from thermal noise. This item continues to be a concern and will be revisited during the Spacecraft CDR.
Systems Engineering activities for the month of January 1996 focused on several design and documentation issues including: ACIS specifications and ICD's update, spacecraft grounding, spacecraft charging, serial digital interfaces between the PSMC and RCTU, test plan/procedures responsibilities, and verification reports responsibilities. Additionally, the Systems Engineering group supported the monthly MIT/LMA TIM held in Denver on 1/25/96 and 1/26/96 and the Spacecraft (TRW) Action Item Telecons.
The ACIS PTS Specification Rev. C, submitted in November of 1995, was released by LMA without MIT signature based on no comments received from MIT. This provides a baseline document for the program. An update to this document will be scheduled for release in the near future to reflect impacts from Change Order #43.
Update of the GSE CEI Specification, planned for completion this month, was delayed as a result Mr. Glen Schaff, lead for MIT EGSE design, leaving MIT employment. Coordination with the new EGSE lead, Demitrios Athens, and information gathering activities were initiated this month to enable update of the GSE CEI Specification for EGSE items by mid-February 1996.
Update of the GSE to ACIS and Facilities ICD is in process. A draft version of this ICD release is scheduled for mid-March 1996 with a final version to be released on or about 4/15/96.
Release of the ACIS PTS to DPS ICD and the ACIS FP to DH ICD was rescheduled to accommodate the MIT/LMA TIM on 1/25/96 and 1/26/96. These documents were reviewed by MIT personnel during the TIM and were subsequently released on 1/28/96 without Change Order #43 impacts. Change Order impacts will be incorporated in the near future.
During this reporting period Systems Design activities continued to support the development and design verification for spacecraft charging requirements by TRW for the AXAF-I mission.
Additionally, the architecture for spacecraft grounding continued to be an issue. During a Spacecraft Grounding TIM held at TRW this month, an issue with respect to RCTU to ACIS (PSMC) Serial Digital Telemetry interfaces was uncovered. This interface issue is planned to be resolved before or during the Spacecraft CDR to be held the first week of February.
The group continued support of the AXAF-I Weekly Action Item Tracking telecon scheduled on Wednesday as well as the weekly MIT/LMA telecon on Thursday.
Update of the ground processing flows for test of the ACIS instrument hardware continued throughout January. LMA continues to maintain the ACIS component and system level test flows as a matter of normal business.
Spacecraft CDR at Huntsville.
Schedule remains the most significant management problem.
Electrical power requirements (Watts) are summarized in the following table:
|Peak power distribution
in Standby Mode
|Peak power distribution
in Max. Operating Mode
|Peak power distribution
in Bakeout Mode
|Peak power distribution in
Normal Operating Mode*
Note: Normal operating mode refers to the ACIS operating with six analog chains at full power, six front-end processors at full power, one back-end processor at full power and the focal plane temperature being maintained at -120 deg C.